OpenStudy (anonymous):

How would you evaluate 4log6(216)

OpenStudy (anonymous):

216 = 6^3 4log6(216) \[^{4}\log 6^{4}\]=\[4^{2^{2}}\log6\] = \[4^{2}\log6^{1/2}\] = \[2 ^{2}\log6\]

OpenStudy (anonymous):

you have to know that \[^{a}logb ^{c} = c ^{a}logb\] \[^{a ^{b}}logc = ^{a}logc ^{(1/b)} = (1/b)^{a}\log c\]

OpenStudy (anonymous):

Okay, thnx

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

df